OpenStudy (anonymous):

17. Model the function rule y = -2x - 3 with a table of values.

OpenStudy (cwrw238):

take some values of x say -3, -2, -1 , 0 . 1. 2. 3 and find the corresponding values of y by substituting them in y = -2x - 3

OpenStudy (anonymous):

okay so how do i find the corresponding values of y?

OpenStudy (anonymous):

@cwrw238

OpenStudy (jim766):

substitute each x in the equation y = -2x -3 for ex y = -2(-3) - 3 simplify the right side and that will get you y for when x = -3

OpenStudy (anonymous):

@Jim766 okay :)

OpenStudy (jim766):

x y -3 3 -2 1 -1 -1 you good

OpenStudy (anonymous):

okay, do i do anything else?

OpenStudy (jim766):

depends how many values you need in the table...but no, that is it you can find as many x's and y's as you want

OpenStudy (anonymous):

okay thank you, could you help me with another question?

OpenStudy (jim766):

sure

OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (jim766):

the x intercept is where y = 0 (____, 0 ) to find this, substitute 0 in for y and solve for x , 0 = -2x-3 the y intercept is where x = 0 ( 0 , ____) y = 2(0) -3

OpenStudy (anonymous):

thank you :)

OpenStudy (jim766):

yw